# Efficient Artifacts Removal for Adaptive Deep Brain Stimulation and a Temporal Event Localization Analysis

### v1

@@ 4.2. Performance evaluation
SMARTA+, SMARTA, template subtraction, and blanking were evaluated using 34 real and semi-real aDBS recording. A quantitative comparison of these methods, alongside template subtraction and blanking, is presented in Table 1 and Table 2 for real signals and semi-real signals, respectively. The accuracy of the estimated artifact-free signals was assessed using NMSE, and artifact removal effectiveness was evaluated using AR in the time domain and SC in the frequency domain. The AR values presented in Table 1 and Table 2 represent the average across all signal segments, while SC values are averaged over specific frequency points: $f_{c}=129.16\times n$ Hz, $42.8+129.16\times m$ Hz, and $86.36+129.16\times m$ Hz for $n=1,2,3$ and $m=0,1,2,3$. The table also lists the average computation time per segment. Both SMARTA and SMARTA+ substantially outperformed template subtraction and blanking methods in all performance metrics, albeit at the cost of higher computational load. Among these two, SMARTA+ performed better than SMARTA in terms of NMSE, AR and SC, indicating that it can more effectively restore the original LFP signal. A significant advantage of SMARTA+ was its reduced computation time, decreasing the average processing time per segment by more than 90% compared to SMARTA. This improvement enhances its feasibility for real-time aDBS applications.
$\mu$ AR SC $\mathrm{T}_{c}$ (ms)
$(\sigma)$
Raw data 2.4892 0.4681 N/A
@@
Transient 0.5560 0.1467 0.0274
blanking (0.1105) (0.1369) (0.0016)
SMARTA 0.1523 0.0789 181.4394
@@
SMARTA+ 0.1216 0.0762 9.3409
(0.0392) (0.0117) (0.6506)
@@
$\mu$ NMSE AR SC $\mathrm{T}_{c}$ (ms)
$(\sigma)$
Raw data 26.7102 2.4688 0.4882 N/A
(6.0883) (0.2806) (0.1499) N/A
Template 18.8519 1.5113 0.4186 0.0157
subtraction (6.3660) (0.4505) (0.1588) (0.0022)
Pulse 14.8655 0.5467 0.4625 0.0249
blanking (5.4402) (0.5073) (0.1042) (0.0070)
Transient 14.9663 0.5581 0.1876 0.0284
blanking (7.1509) (0.1116) (0.1725) (0.0020)
SMARTA -0.8739 0.0738 0.0612 83.9569
(5.2589) (0.0146) (0.0048) (1.3635)
SMARTA+ -10.6868 0.0652 0.0602 7.9830
(2.0356) (0.0089) (0.0055) (0.2417)
Table 2. Performance comparison of artifact removal methods on semi-real aDBS signals. $\mu$: mean; $\sigma$: standard deviation; NMSE: normalized mean square error; $\mathrm{NMSE}_{\beta}$: NMSE of $\beta$ band (13-35Hz); AR: artifact residual; SCSC: spectral concentration; $\mathrm{T}_{c}$: computation time per segment; N/A: not applicable.
To further examine artifact removal performance across frequency bands, NMSE values were computed for alpha (4-8 Hz), beta (13-35 Hz), gamma (60-90 Hz), high-frequency oscillation (HFO; 200-400 Hz), and very high-frequency oscillation (VHFO; 400-1000 Hz and 1000-3000 Hz for VHFO1 and VHFO2, respectively) bands. The results are summarized in Table 3. These values were obtained by applying band-pass filters to the signals before NMSE calculation. In the semi-real aDBS signals, the artifact signal was generated separately by averaging over 500 segments before added to the base LFP signal. This extensive averaging effectively removed most of the underlying beta-band activity from the artifacts, resulting in minimal distortion of beta activity in the raw semi-real data, as reflected by the low beta-band NMSE. The alpha band exhibited high NMSE values in raw signals due to low-frequency drift caused by DC transient artifacts. Only SMARTA+, which incorporates a line-fitting algorithm for transient correction, demonstrated a significant reduction in NMSE in this band. In the gamma band, the aliasing of stimulation harmonics, particularly at 86.36 Hz, contributed to elevated NMSE in raw data. SMARTA+ again yielded the lowest NMSE, outperforming other compared algorithms. The HFO and VHFO bands contained a broader spectrum of artifact-induced distortion, including both harmonics and aliasing. As expected, the NMSE values for the raw semi-real data in this band were the highest. While transient blanking significantly improved HFO-band and VHFO-band NMSE compared to pulse blanking and template subtraction, SMARTA+ achieved the most substantial reduction among all methods. This suggests its superior capacity to suppress both time-domain and frequency-domain components of stimulation artifacts. Overall, SMARTA+ outperformed all other methods across all four frequency bands, supporting its effectiveness in restoring LFP signals over the full spectral range.
$\mu$ $\alpha$ $\beta$ $\gamma$ HFO VHFO1 VHFO2
$(\sigma)$
Raw data 8.8129 -2.8386 6.5825 22.0838 22.0623 28.3007
@@
SMARTA+ -5.8872 -13.5504 -6.9826 -10.3981 -12.1729 -12.4535
(2.0356) (4.4098) (3.5192) (4.9695) (2.9736) 2.3763
@@ 4.3. Temporal Event Localization Analysis
Table 4 presents the accuracy of the beta event detection algorithm, using the baseline beta amplitude extraction procedure described in Sec. 2.1, following the application of various artifact removal methods. Below, “Raw filtering” denotes the condition in which no artifact suppression techniques was applied before evaluating beta amplitude. Transient blanking [1] was intended to avoid artifact-induced self-triggering by suppressing beta detection for a fixed duration following stimulation onset. However, because it simply omitted the first 550 ms of each stimulation period, it did not contribute to accurate beta event detection and could not be meaningfully assessed. As such, transient blanking was excluded from direct performance comparisons.
@@
There are several findings. First, when no artifact removal was applied; that is, raw-filtering, the extracted beta amplitude resulted in high recall but low precision. This pattern was especially pronounced during the onset of stimulation, suggesting that many of the detected beta events were in fact false positives caused by artifacts overlapping with the beta band. The inflated recall, therefore, reflected the detection algorithm’s susceptibility to spurious high-amplitude signals, rather than its ability to identify true neural activity. Second, template subtraction, pulse blanking, and SMARTA all demonstrated improvements upon raw-filtering in detection accuracy by reducing the number of false positives. These method led to a notable increase in precision, thereby improving the F1-score and reducing timing deviation. Among them, SMARTA+ achieved the highest overall performance, particularly at the onset of stimulation. It demonstrated significantly higher precision and F1-score than other approaches, while maintaining a comparable recall. In addition, SMARTA+ notably reduced the timing deviation and increased the overlap ratio in beta event detection, providing more temporally precise alignment between the detected beta activity and the underlying neural signal.
@@
SMARTA+ 91.12 92.18 92.81 93.25 91.81 92.63 59.20 38.65 92.62 94.07 88.72 89.71
Table 4. Temporal event localization accuracy comparison of the aDBS algorithm before and after applying artifact removal methods. The p-value obtained using the paired Wilcoxon signed-rank test between SMARTA+ and another algorithm is shown in the parenthesis, where $\ast$ means $p<10^{-3}$.

### v2

@@ 4.2. Performance evaluation
SMARTA+, SMARTA, template subtraction, and blanking were evaluated using 34 real and semi-real aDBS recordings. A quantitative comparison of these methods, alongside template subtraction and blanking, is presented in Table 1 and Table 2 for semi-real signals and real signals, respectively. The accuracy of the estimated artifact-free signals was assessed using NMSE, and artifact removal effectiveness was evaluated using AR in the time domain and SC in the frequency domain. The AR values presented in Table 1 and Table 2 represent the median across all signal segments, while SC values are averaged over specific frequency points: $f_{c}=129.16\times n$ Hz, $42.8+129.16\times m$ Hz, and $86.36+129.16\times m$ Hz for $n=1,2,3$ and $m=0,1,2,3$. The table also lists the average computation time per segment, denoted as $T_{c}$. Recall that due to the construction, SMARTA+ achieved stable performance after a few stimulation periods (See Figure 4). To ensure a fair evaluation of the model’s effective performance, in addition to results over the entire recording, we also report performance metrics with the first three stimulation periods excluded. The corresponding results are presented in the last row of the tables.
For semi-real signals, as shown in Table 1, both SMARTA and SMARTA+ substantially outperformed the template subtraction and blanking methods across all metrics, except for SC values of SMARTA+ when the first three stimulation periods were included. Between the two, SMARTA+ yielded lower NMSE and AR value but slightly inferior SC value relative to SMARTA, suggesting comparable capability in restoring the underlying LFP signals. For real aDBS signals, as shown in Table 2, consistent trends were observed. A significant advantage of SMARTA+ was its substantially lower computation time, decreasing the average processing time per segment by more than 90% compared to SMARTA. This improvement enhances its feasibility for real-time aDBS applications. The average time required to rebuild the decision trees in SMARTA+ was $346.24\pm 58.41$ ms, in addition to the artifact-removal computation time. The mean non-stimulation interval between successive stimulation periods was $588.08\pm 305.75$ ms, which allowed the decision-forest rebuilding procedure to be executed after each stimulation period and completed before the next.
Median NMSE AR SC $\mathrm{T}_{c}$ (ms)
MAD
Raw data 25.0115 2.5192 0.4376 N/A
(4.0334) (0.2237) (0.0938) N/A
Template 19.2288 1.5660 0.3791 0.0476
subtraction (5.1384) (0.3218) (0.1106) (0.0012)
Pulse 12.5190 0.1372 0.4446 0.0947
blanking (2.9617) (0.1077) (0.0558) (0.0014)
Transient 14.0001 0.4670 0.1369 0.1015
blanking (5.3030) ($10^{-4}$) (0.0810) (0.0029)
SMARTA -0.6043 0.0393 0.0577 83.7437
(3.4507) (0.0033) (0.0044) (1.1043)
SMARTA+ -1.6195 0.0343 0.1814 2.2379
(whole signal) (3.3831) (0.0033) (0.0565) (0.0872)
SMARTA+ -9.0256 0.0333 0.0633 2.2439
(stable periods) (7.7040) (0.0030) (0.0072) (0.0816)
Table 1. Performance comparison of artifact removal methods on semi-real aDBS signals. MAD: median absolute deviation; NMSE: normalized mean square error; $\mathrm{NMSE}_{\beta}$: NMSE of $\beta$ band (13-35Hz); AR: artifact residual; SCSC: spectral concentration; $\mathrm{T}_{c}$: computation time per segment; N/A: not applicable; stable periods: excluding the first three stimulation periods for evaluation.
Median AR SC $\mathrm{T}_{c}$ (ms)
Raw data 2.5175 0.4346 N/A
@@
Transient 0.4670 0.1354 0.0958
blanking ($10^{-4}$) (0.0791) (0.0008)
SMARTA 0.0625 0.0894 83.0630
@@
SMARTA+ 0.0530 0.1892 3.5924
(whole signal) (0.0131) (0.0516) (0.1127)
SMARTA+ 0.0515 0.1073 3.5951
(stable periods) (0.0129) (0.0225) (0.0961)
@@
4.3. Different spectral bands
To further examine artifact removal performance across frequency bands, NMSE values were computed for alpha (4-8 Hz), beta (13-35 Hz), gamma (60-90 Hz), high-frequency oscillation (HFO; 200-400 Hz), and very high-frequency oscillation (VHFO; 400-1000 Hz and 1000-3000 Hz for VHFO1 and VHFO2, respectively) bands. The results are summarized in Table 3. These values were obtained by applying band-pass filters to the signals before NMSE calculation. In the semi-real aDBS signals, the artifact signal was generated separately by averaging over 500 segments before added to the base LFP signal. This extensive averaging effectively removed most of the underlying beta-band activity from the artifacts, resulting in minimal distortion of beta activity in the raw semi-real data, as reflected by the low beta-band NMSE. The alpha band exhibited high NMSE values in raw signals due to low-frequency drift caused by DC transient artifacts. Only SMARTA+, which incorporates a line-fitting algorithm for transient correction, demonstrated a significant reduction in NMSE in this band. In the gamma band, the aliasing of stimulation harmonics, particularly at 86.36 Hz, contributed to elevated NMSE in raw data. SMARTA+ again yielded the lowest NMSE, outperforming other compared algorithms. The HFO and VHFO bands contained a broader spectrum of artifact-induced distortion, including both harmonics and aliasing. As expected, the NMSE values for the raw semi-real data in the HFO and VHFO bands were the highest. While transient blanking substantially improved NMSE in the HFO and VHFO bands compared with pulse blanking and template subtraction, SMARTA achieved the greatest overall reduction among these methods. When considering the stable periods of SMARTA+, its improvement in the HFO band exceeded that of SMARTA, while its performance in the VHFO band remained comparable. These findings confirm the superior performance of SMARTA+ in suppressing stimulation artifacts across time and frequency domains, yielding the most accurate restoration of LFP signals over the entire spectral range.
Med $\alpha$ $\beta$ $\gamma$ HFO VHFO1 VHFO2
Raw data 9.2380 -3.4122 7.1040 21.9413 21.6738 26.6248
@@ 4.3. Different spectral bands
SMARTA+ -5.4408 -13.5463 -5.3779 -1.9351 -3.0674 -2.0688
(whole signal) (1.4129) (3.1985) (2.8050) (6.0309) (4.9093) (4.0505)
SMARTA+ -5.8774 -13.5198 -5.8394 -9.1458 -10.2569 -10.3131
(stable periods) (1.3845) (3.2608) (2.8709) (3.6409) (2.0767) (1.4184)
@@ 4.4. Temporal Event Localization Analysis
Table 4 presents the accuracy of the beta event detection algorithm, using the baseline beta amplitude extraction procedure described in Sec. 2.1, following the application of various artifact removal methods. Below, “Raw filtering” denotes the condition in which no artifact suppression techniques was applied before evaluating beta amplitude. Transient blanking [anso2022artifactRemoval] was intended to avoid artifact-induced self-triggering by suppressing beta detection for a fixed duration following stimulation onset. However, because it simply omitted the first 550 ms of each stimulation period, it did not contribute to accurate beta event detection and could not be meaningfully assessed. As such, transient blanking was excluded from direct performance comparisons.
@@
There are several findings. First, when no artifact removal was applied; that is, raw-filtering, the extracted beta amplitude resulted in high recall but low precision. This pattern was especially pronounced during the onset of stimulation, suggesting that many of the detected beta events were actually false positives caused by artifacts overlapping with the beta band. The inflated recall, therefore, reflected the detection algorithm’s susceptibility to spurious high-amplitude signals, rather than its ability to identify true neural activity. Second, template subtraction, pulse blanking, and SMARTA all demonstrated improvements upon raw-filtering in detection accuracy by reducing the number of false positives. These method led to a notable increase in precision, thereby improving the F1-score and reducing timing deviation. Among them, SMARTA+ achieved the highest overall performance, particularly at the onset of stimulation. It demonstrated significantly higher precision and F1-score than other approaches, while maintaining a comparable recall. In addition, SMARTA+ notably reduced the timing deviation and increased the overlap ratio in beta event detection, providing more temporally precise alignment between the detected beta activity and the underlying neural signal.
The performance of SMARTA+ evaluated on the stable periods exceeded that obtained using the entire signal, suggesting that incomplete artifact suppression during the initial stimulation periods may introduce minor inaccuracies in beta-event detection. For most metrics, however, this difference was not statistically significant, and beta events were recovered with accuracy comparable to that observed during the stable phase.
@@
(0.80) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$)
SMARTA 92.73 91.83 63.10 78.82 73.21 84.50 163.59 95.34 74.09 81.66 88.27 90.48
(0.73) (0.004) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) (0.04) ($\ast$)
SMARTA+ 92.54 93.30 90.96 90.77 91.63 91.91 55.58 36.14 91.44 92.19 90.77 92.24
(whole signal) (0.02) (0.27) (0.26) (1.00) (0.35) (0.82) (0.48) (0.03) (0.59) (0.10) (0.06) (0.45)
SMARTA+ 92.48 93.29 91.14 90.80 91.69 91.92 54.67 35.68 91.62 92.25 90.58 92.31
(stable periods)
Table 4. Temporal event localization accuracy comparison of the aDBS algorithm before and after applying artifact removal methods. The p-value obtained using the paired Wilcoxon signed-rank test between SMARTA+ (stable periods) and another algorithm is shown in the parenthesis, where $\ast$ means $p<10^{-3}$.
4.5. Low sampling rate
Table 5 presents the results obtained from semi-real signals resampled to 1.1 kHz, aiming to evaluate the performance of SMARTA+ on data with a sampling rate closer to that used in clinical DBS systems. Although the NMSE and SC scores of SMARTA and SMARTA+ showed slight degradation, likely due to distortion introduced during resampling, both methods consistently outperformed template subtraction and blanking. For the AR metric, only three and six points were available for computing $\hat{Z}$ and $\hat{Z}$ in Eq. 3, respectively, which biased the median estimate. Consequently, the first term in Eq. 3 became large, suggesting that AR may not be an appropriate measure for evaluating low-sampling-rate signals. Overall, these results indicate that SMARTA+ remains effective even for signals acquired at reduced sampling rates.
Table 5 also summarizes the NMSE across different frequency bands. Note that VHFO components could not be evaluated for signals sampled at 1.1 kHz. Overall, the results were consistent with those in Table 3, with the exception that transient blanking achieved the lowest NMSE in the $\gamma$ band. This likely reflects the fact that transient blanking is less affected by resampling-induced distortion compared with template-based methods.
The temporal event localization results for the 1.1 kHz signals are presented in Table 6. Although overall performance decreased, the findings remained consistent with those in Table 4.
Median NMSE AR SC $\mathrm{T}_{c}$ (ms) $\alpha$ $\beta$ $\gamma$ HFO
Raw data 15.8987 1.1925 0.4378 N/A 9.2494 -3.4084 7.1020 22.0712
(3.6016) (0.4462) (0.0940) N/A (3.4453) (4.4811) (3.7120) (3.7425)
Template 11.0212 0.7166 0.2220 0.0299 9.4507 -3.8814 7.7599 16.4511
subtraction (3.7542) (0.2589) (0.0536) (0.0011) (3.6498) (4.6678) (4.6230) (4.8424)
Pulse 15.9391 0.8938 0.3290 0.1001 9.6025 -2.5491 14.8643 21.1013
blanking (3.6495) (0.2144) (0.0345) (0.0024) (3.0235) (4.0273) (3.6778) (3.7700)
Transient 7.7286 0.2231 0.1634 0.1021 3.8334 -3.2304 1.5065 5.1021
blanking (5.1407) ($10^{-4}$) (0.0941) (0.0032) (3.3030) (1.5004) (5.2869) (6.8343)
SMARTA 0.5787 0.5157 0.1175 34.9759 9.0866 -4.1982 3.0470 -0.6028
(4.1327) (0.0171) (0.1175) (0.2794) (3.2960) (3.6839) (3.2201) (2.3685)
SMARTA+ -2.2357 0.5332 0.1808 2.9273 -4.6041 -10.1709 2.9351 3.6570
(whole signal) (3.4422) (0.0255) (0.0462) (0.1810) (1.4740) (4.2665) (2.8059) (4.1167)
SMARTA+ -3.5628 0.5343 0.1547 2.9042 -5.2826 -10.3744 2.1548 2.3149
(stable periods) (3.0709) (0.0219) (0.0447) (0.1761) (1.5635) (4.1279) (3.0291) (3.6215)
Table 5. Performance comparison of artifact removal methods on semi-real aDBS signals. The last four columns show the band-specific NMSE values. All signals were resampled to 1.1 kHz.
Recall (%) Precision (%) F1-score (%) deviation (ms) OR1 (%) OR2 (%)
Onset All Onset All Onset All Onset All Onset All Onset All
Raw filtering 91.78 91.80 58.38 77.30 69.28 83.61 172.51 101.42 71.64 81.13 87.86 89.71
@@ 4.5. Low sampling rate
SMARTA+ 90.45 92.15 86.13 86.42 87.97 89.02 80.10 48.84 87.23 89.50 89.39 90.95
(whole signal) (0.005) (0.94) (0.46) (0.01) (0.30) (0.04) (0.50) (0.007) (0.21) (0.002) (0.49) (0.51)
SMARTA+ 90.17 92.18 86.44 86.63 87.95 89.14 78.64 48.12 87.62 89.63 89.40 91.02
(stable periods)
Table 6. Temporal event localization accuracy comparison of the aDBS algorithm before and after applying artifact removal methods. All signals were resampled to 1.1 kHz.
4.6. Different stimulation frequency
To evaluate the flexibility of SMARTA+ under different stimulation settings, semi-real aDBS signals with a stimulation frequency of 180 Hz were generated and used for artifact-removal assessment. The total number of reconstructed artifacts per signal was matched to the 130 Hz condition, while the inter-pulse interval was adjusted to 1/180 s. As a result, the stimulation periods in this dataset were shorter than those in the original aDBS recordings and in the semi-real signals at 130 Hz. Table 7 presents the performance scores for all methods tested on the 180 Hz signals. The results were consistent with those in Table 4, with the exception of the AR value for transient blanking. This difference arises because the blanking duration was fixed at 550 ms; thus, due to the shorter stimulation periods at 180 Hz, most stimulation segments were blanked out. Consequently, the AR index was computed largely over flat segments, resulting in values approaching zero.
The findings for the NMSE scores across frequency bands and the temporal localization results for the 180-Hz signals, which is presented in Table 8, are consistent with those observed for the 130-Hz semi-real signals.
Median NMSE AR SC $\mathrm{T}_{c}$ (ms) $\alpha$ $\beta$ $\gamma$ HFO VHFO1 VHFO2
MAD
Raw data 22.4071 2.9329 0.1220 N/A 9.5436 -0.9062 4.9099 12.8157 12.4137 17.0511
(5.9845) (0.1845) (0.0067) N/A (3.9991) (4.4513) (3.0578) (3.3711) (3.2385) (3.6469)
Template 17.9437 2.9593 0.1758 0.0402 8.5478 -2.1402 4.7517 9.0439 9.5489 14.8217
subtraction (5.8616) (0.2424) (0.0185) (0.0008) (4.0624) (4.5610) (2.7667) (3.4116) (3.1729) (4.0371)
Pulse 9.7193 0.3279 0.1789 0.0916 8.9567 -1.2660 6.8580 11.6775 9.1690 11.4547
blanking (3.4355) (0.1497) (0.0422) (0.0015) (4.3540) (4.4017) (3.8403) (3.8889) (3.3216) (3.8824)
Transient 7.3404 0.0004 0.1636 0.0979 5.5571 -1.7524 0.0221 6.9245 7.1249 8.4059
blanking (2.9605) ($10^{-4}$) (0.0257) (0.0025) (2.8232) (2.3255) (2.9927) (3.6900) (3.4426) (3.9062)
SMARTA 2.7747 0.0547 0.2061 74.8882 8.3759 -1.8588 0.3693 -0.5113 -9.6180 -11.4659
(3.8402) (0.0047) (0.0068) (0.7511) (4.4234) (4.8607) (2.7084) (3.0941) (2.3253) (2.5624)
SMARTA+ -3.2299 0.0436 0.0592 2.0086 -4.9586 -10.9213 -1.9033 -5.5759 -7.6421 -8.1893
(whole signal) (4.2170) (0.0031) (0.0095) (0.0760) (1.3371) (3.5000) (3.1899) (3.6052) (3.1344) (2.6886)
SMARTA+ -8.5191 0.0422 0.0555 2.0130 -5.0670 -11.0432 -2.6836 -7.9150 -9.7905 -10.9005
(stable periods) (2.4420) (0.0027) (0.0056) (0.0688) (1.5979) (3.7163) (3.2190) (2.8310) (2.4554) (1.6522)
Table 7. Performance comparison of artifact removal methods on semi-real aDBS signals with a stimulation frequency of 180 Hz. The last six columns show band-specific NMSE values.
Recall (%) Precision (%) F1-score (%) deviation (ms) OR1 (%) OR2 (%)
Onset All Onset All Onset All Onset All Onset All Onset All
Raw filtering 93.24 91.46 63.09 77.58 73.01 83.51 214.21 109.35 71.98 82.17 87.06 91.15
Template 92.85 92.13 63.89 77.42 73.35 83.68 202.08 104.07 72.68 82.18 86.91 90.63
subtraction (0.19) (0.03) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) (0.03)
Blanking 93.76 91.88 69.27 78.46 77.73 84.10 196.61 103.20 74.18 82.62 87.66 90.96
SMARTA 92.77 92.25 66.32 78.39 75.18 84.39 183.48 99.04 74.53 82.70 88.22 90.68
SMARTA+ 94.38 93.63 91.84 90.29 92.99 91.86 61.44 38.68 91.28 92.45 90.72 92.01
(whole signal) (0.31) (0.20) (0.91) (0.60) (0.52) (0.60) (0.17) ($\ast$) (0.63) (0.005) (0.63) (0.06)
SMARTA+ 94.58 93.62 92.40 90.30 93.35 91.87 60.29 37.91 91.39 92.60 90.82 92.17
(stable periods)
Table 8. Temporal event localization accuracy comparison of the aDBS algorithm before and after applying artifact removal methods on semi-real signals with a stimulation frequency of 180 Hz.